OpenStudy (anonymous):

2x+square root of x -6=0?

OpenStudy (anonymous):

let us assume sqrtx = y therefore x = y^2 now plug in these values in above equation and solve for y

OpenStudy (anonymous):

alright. could you finish just i understand exactly?

OpenStudy (anonymous):

\[2y^2 +y -6 =0\] now solve for y

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

2y^2+y=-6

OpenStudy (anonymous):

then...?

OpenStudy (anonymous):

got lost :/

OpenStudy (anonymous):

\[2y^2 + y - 6=0\] \[2y^2 + 4y -3y - 6=0\] \[2y(y + 2) -3(y + 2)=0\] \[(y + 2)(2y -3)=0\] therefore either y + 2 = 0 or, 2y-3 = 0 therefore y = -2 or, y = 3/2 now plug in these value in y = sqrtx

OpenStudy (anonymous):

ty ty ty